# CHOSEN: Contrastive Hypothesis Selection for Multi-View Depth Refinement

### v1

@@ 4.2 Ablation studies
  • ($\nicefrac{{1}}{{8}}$,$\nicefrac{{1}}{{4}}$,$\nicefrac{{1}}{{4}}$)-($\nicefrac{{1}}{{4}}$,$\nicefrac{{1}}{{2}}$,$1$): There are 3 stages of refinement with $\nicefrac{{1}}{{8}}\,\&\,\nicefrac{{1}}{{4}}$, $\nicefrac{{1}}{{4}}\,\&\,\nicefrac{{1}}{{2}}$ and 1/8& $\nicefrac{{1}}{{8}}\,\&\,1$ cost volume pyramids, and output refined depth at $\nicefrac{{1}}{{4}}$ resolution. This is our default configuration and we also include the results trained on data mixed BlendedMVS.
@@
The ”tamed” second order error term $\mathbf{e}(d)$ in Eq.(4) can be viewed as an extra component compared to the input feature design in the RAFT (Teed & Deng, 2020) framework for optical flow estimation. Here we show that it is a vital component in the hypothesis ranking model that improves the overall smoothness and accuracy. Quantitative results on DTU testing set are reported in the sixth row of Tab.1. Since each hypothesis is evaluated independently, without the information about how well a particular hypothesis fits in the local geometry, the model struggles to select the best hypothesis solely based on matching and context information.
@@ 4.3 Comparisons on depths and normals
4.4 Qualitative results
We demonstrate the excellent generalization ability of our simple baseline on various datasets including MultiFace (Wuu et al., 2022), Tanks & Temples (Knapitsch et al., 2017) and ETH3D (Schops et al., 2017). Results in Fig.4 use the same baseline model trained on mixture of DTU and BlendedMVS reported in Tab.1.
@@ 5 Limitations and Conclusion
One limitation of CHOSEN is that it becomes more expensive to sample and select from a large number of hypotheses at high resolution. This is part of the reason why we settled at $\nicefrac{{1}}{{4}}$ resolution to arrive at a sweet spot of both performance and run time. Another limitation is that we did not focus on bundle adjustment and joint filtering of depths for point cloud and surface reconstruction, which we believe should be a subject of its own interests, especially considering the proliferation of volumetric based methods such as (Zhang et al., 2021).
We have demonstrated that CHOSEN is simple yet effective for multi-view depth refinement. It iteratively re-samples and select the best hypotheses, and is applicable to a wide range of multi-view capture setup. We have integrated CHOSEN into a simple MVS baseline that delivers impressive quality in terms of depth and normal accuracy.

### v2

@@ 4.2 Ablation studies
  • ($\nicefrac{{1}}{{8}}$,$\nicefrac{{1}}{{4}}$,$\nicefrac{{1}}{{4}}$)-($\nicefrac{{1}}{{4}}$,$\nicefrac{{1}}{{2}}$,$1$): There are 3 stages of refinement with $\nicefrac{{1}}{{8}}\,\&\,\nicefrac{{1}}{{4}}$, $\nicefrac{{1}}{{4}}\,\&\,\nicefrac{{1}}{{2}}$ and 1/4& $\nicefrac{{1}}{{4}}\,\&\,1$ cost volume pyramids, and output refined depth at $\nicefrac{{1}}{{4}}$ resolution. This is our default configuration and we also include the results trained on data mixed BlendedMVS.
@@
The ”tamed” second order error term $\mathbf{e}(d)$ in Eq.(4) can be viewed as an extra component compared to the input feature design in the RAFT [38] framework for optical flow estimation. Here we show that it is a vital component in the hypothesis ranking model that improves the overall smoothness and accuracy. Quantitative results on DTU testing set are reported in the sixth row of Tab.1. Since each hypothesis is evaluated independently, without the information about how well a particular hypothesis fits in the local geometry, the model struggles to select the best hypothesis solely based on matching and context information.
@@ 4.3 Comparisons on depths and normals
4.4 Qualitative results for baseline
We demonstrate the excellent generalization ability of our simple baseline on various datasets including MultiFace [42], Tanks & Temples [24] and ETH3D [35]. Results in Fig.5 use the same baseline model trained on mixture of DTU and BlendedMVS reported in Tab.1.
4.5 A simple extension of the baseline MVS with cost volume filtering
We explore a very simple extended setting to our baseline MVS, where we apply an MVSNet [44] style 3D-CNN on our full range cost volume, while other components remains unchanged. The 3D-CNN is applied on the 1-channel correlation volume computed according to Eq.(10). The 3D-CNN then outputs the logit for each depth hypothesis, and the logits per pixel are further normalized by the soft-max function. This provides more global matching information to CHOSEN.
Compared to the correlation volume in the baseline, the 3D-CNN and the normalization allow CHOSEN benefit from the higher quality matching signal. Thus we can use just $\mathcal{L}_{\text{cl}}$ loss to supervise the whole pipeline, without the loss $\mathcal{L}_{\text{fm}}$ in our baseline MVS. We found that this model produces performs much better on inputs with large texture-less regions, such as from the ETH3D datasets shown in Figure 6. We thus believe our formulation of CHOSEN will present mutual benefit with many of the current MVS frameworks that focus on cost volume and feature quality.
@@ 5 Limitations and Conclusion
We have demonstrated that CHOSEN is simple yet effective for multi-view depth refinement. One limitation of CHOSEN is that it becomes more expensive to sample and select from a large number of hypotheses at high resolution. This is part of the reason why we settled at $\nicefrac{{1}}{{4}}$ resolution to arrive at a sweet spot of both performance and run time. Another limitation is that we did not focus on bundle adjustment and joint filtering of depths for point cloud and surface reconstruction, which we believe should be a subject of its own interests, especially considering the proliferation of volumetric based methods such as [48].
Supplementary Material
Motivation for the baseline design
Our baseline MVS design has at least the following disadvantages:
  • the matching features do not receive gradients from the hypothesis selection module.
  • simple feature cross correlation is used, which is very noisy.
  • there is no cost volume processing to denoise the matching cost volume
This matching component is thus isolated from the selection module and has sub-optimal performance, struggling in case of large texture-less areas, perspective distortions, etc. An example evaluation result of the matching component only is given in Tab.1, row titled “baseline w/o CHOSEN”. Even with the noisy matching cost and its noisy initial depth, the selection module is still powerful enough to yield high quality depth estimation that surpasses other state-of-art models in terms of depth and normal accuracy on DTU. Therefore, one should expect better performance if we have a cost volume that behaves better than our noisy cross correlation. We did NOT aim to select a canonical architecture for matching cost, which is the focus of many other frameworks. We did demonstrate that CHOSEN can recover good estimation even if the matching cost is very noisy.
Architecture details
We describe the details of the U-Nets used in the baseline MVS in Tab.2. A block refers to a sequence of convolution-ReLU-batch_norm layers, with specified kernel size and dilation rate. The context features $\mathbf{c}$ for CHOSEN contain 64 and 48 channels at $\nicefrac{{1}}{{8}}$ and $\nicefrac{{1}}{{4}}$ resolutions. These features will be reinitialized at a new cost volume config level with two 3x3-convolution layers. When updating them during the iteration, we use two ConvGRUs, one operate at resized input features to half of the size and one operate at full-size input features, and combine them through addition. The ranking module has four 1x1-convolution layers, with intermediate channel size 64. Our baseline MVS model is very small in terms of parameter count compared to other frameworks.
U-Net: matching and context feature extractors
Block name Kernel sizes Dilations strides Channels I/O In res. Out res. Input name
unet_conv0 [4, 3] [1, 1] [1, 1] 3/16 1 1 $I_{\rm RGB}$
unet_conv1 [4, 3] [1, 1] [2, 1] 16/32 1 1/2 unet_conv0
unet_conv2 [4, 3] [1, 1] [2, 1] 32/48 1/2 1/4 unet_conv1
unet_conv3 [4, 3] [1, 1] [2, 1] 48/48 1/4 1/8 unet_conv2
unet_conv4* [4, 3] [1, 1] [2, 1] 48/48 1/8 1/16 unet_conv3
unet_conv5* [3, 3] [1, 1] [2, 1] 96/48 1/16 1/8 unet_conv3, unet_conv4
unet_conv6* [3, 3] [1, 1] [2, 1] 96/48 1/8 1/4 unet_conv2, unet_conv5
unet_conv7* [3, 3] [1, 1] [2, 1] 80/32 1/4 1/2 unet_conv1, unet_conv6
unet_conv8* [3, 3] [1, 1] [2, 1] 48/16 1/2 1 unet_conv0, unet_conv7
Table 2: Network Architecture details for the matching and context U-Nets in the baseline MVS.
We explore a very simple extended setting to our baseline MVS, where we apply an MVSNet [44] style 3D-CNN on our full range cost volume, while other components remains unchanged. The 3D-CNN is applied on the 1-channel correlation volume computed according to Eq.(10). The 3D-CNN then outputs the logit for each depth hypothesis, and the logits per pixel are further normalized by the soft-max function. This provides more global matching information to CHOSEN.
Compared to the correlation volume in the baseline, the 3D-CNN and the normalization allow CHOSEN benefit from the higher quality matching signal. Thus we can use just $\mathcal{L}_{\text{cl}}$ loss to supervise the whole pipeline, without the loss $\mathcal{L}_{\text{fm}}$ in our baseline MVS. We found that this model produces performs much better on inputs with large texture-less regions, such as from the ETH3D datasets shown in Figure 6. We thus believe our formulation of CHOSEN will present mutual benefit with many of the current MVS frameworks that focus on cost volume and feature quality.
Additional results on MultiFace
We found MultiFace dataset a good testing ground (good calibration, high quality images) and close to the application scenarios our method is originally designed for. In Fig. we provide more qualitative results. Note that the meshes provided in MultiFace dataset are too coarse to be ground truth for meaningful quantitative evaluation.
Figure 7: Direct application of our DTU [1] + BlendedMVS [46] trained baseline model on instances from MultiFace [42].
Figure 8: Direct application of our DTU [1] + BlendedMVS [46] trained baseline model on instances from MultiFace [42], continued from Fig.7